OpenStudy (anonymous):

how to find the derivative of y=e^cos(sqrt3x)?

OpenStudy (anonymous):

We have: \[{y=e^{\cos(\sqrt{3x})}}\] Right?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

Then we can proceed: \[\frac{d}{dx}e^{\cos(\sqrt{3x})}=\frac{d}{dx}e^{\cos\left(\sqrt{f(x)}\right)}=\frac{d}{dx}e^{\cos[g(f(x))]}=\frac{d}{dx}e^{h[g(f(x))]}=\frac{d}{dx}j\bigg({h\Big[g[f(x)]\Big]}\bigg)\] Using the chain rule, we can simplify this

OpenStudy (anonymous):

Make Sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OKAY! SO lets start. Through that little simplification above, we obtained two pieces of information: One is that: \[\frac{d}{dx}e^{\cos(\sqrt{3x})}=\frac{d}{dx}j\bigg({h\Big[g[f(x)]\Big]}\bigg)\] And the second is a result of the first: \[\eqalign{ &f(x)=3x \\ &g(x)=\sqrt{x} \\ &h(x)=cos(x) \\ &j(x)=e^x \\ }\] So then we can use the chain rule to figure out: \[\frac{d}{dx}j\bigg({h\Big[g[f(x)]\Big]}\bigg)=j\phantom{.}'\bigg({h\Big[g[f(x)]\Big]}\bigg)\times h'\Big[g[f(x)]\Big]\times g'[f(x)]\times f'(x)\] So let us find the derivatives for each! \[\eqalign{ &f(x)=3x\phantom{spc}\rightarrow f'(x)=3\\ &g(x)=\sqrt{x}\phantom{spc}\rightarrow g'(x)=\frac{1}{2\sqrt{x}}\\ &h(x)=cos(x)\phantom{spc}\rightarrow h'(x)=-sin(x)\\ &j(x)=e^x\phantom{spc}\rightarrow j'(x)=e^x\\ }\] Now it's just a little algebra. Kinda straightforward, but kinda messy too haha

OpenStudy (anonymous):

\[\eqalign{ &j\phantom{.}'\bigg({h\Big[g[f(x)]\Big]}\bigg)\times h'\Big[g[f(x)]\Big]\times g'[f(x)]\times f'(x) \\ =&e^{\cos(\sqrt{3x})}\times -\sin(\sqrt{3x})\times\frac{1}{2\sqrt{3x}}\times3 \\ =&\frac{-3e^{\cos(\sqrt{3x})}\sin\sqrt{3x}}{2\sqrt{3x}}\\ }\]

OpenStudy (anonymous):

is that the answer??

OpenStudy (isaiah.feynman):

Sorry but his solution is too complicated. Look at the derivative of the exponential function. \[(e^{x} )\prime = e^{x} \times x \prime = e^{x} \times 1 =e^{x}\]

OpenStudy (anonymous):

Yes that's the final answer

OpenStudy (isaiah.feynman):

@KeithAfasCalcLover your solution is complex but your answer is right.

OpenStudy (anonymous):

The function is complex in the sense that it is the composite of four functions and so If the function is complex, so will the answer to give a THOUGHROUGH answer.

OpenStudy (isaiah.feynman):

Wow you seem to have a deep understanding of this. I didn't think of it from that point of view.

OpenStudy (anonymous):

Is that irony? Besides, my answer wasn't even that complex. Just emphasizing the chain rule with more than two chain links. And I thought it was pretty thorough...

OpenStudy (isaiah.feynman):

Usually when I take derivatives of exponential functions I just use the basic exponential function as a reference.

OpenStudy (anonymous):

And what if they have a inner function?

OpenStudy (isaiah.feynman):

Resort to decomposing. lol